Given the following bond energies and the heat of reaction for the process below, estimate the C-H bond energ!y. Bond BDE (kJ/mol) 348 614 436 A. 420 kJ B. 283 kJ C. 525 KJ D. 350 kJ E. 645 kJ
Answer #1

delta H = BE(reactant) - BE(product)

-137 = BE(C=C) + 4*BE(C-H) + BE(H-H) - BE(C-C) - 6*BE(C-H)

-137 = BE(C=C) + BE(H-H) - BE(C-C) - 2*BE(C-H)

-137 = 614 + 436 - 348 - 2*BE(C-H)

2*BE(C-H) = 839

BE(C-H) = 420 KJ

Answer: A